# Corresponding author

### v1

@@ 1 Introduction
The field of Autonomous Mobile Robots (AMRs) is experiencing a surge of development across diverse industries such as healthcare, security monitoring [1][2], and construction [3]. To thrive in these dynamic environments and fulfill their designated tasks, AMRs necessitate robust capabilities in three critical areas: information gathering [4], environment perception [5], and precise Simultaneous Localization and Mapping (SLAM) [6]. Although various research efforts have explored effective mapping techniques that utilize individual sensors such as LiDAR [7] or cameras [8], cameras have significant limitations. Their sensitivity to light fluctuations can considerably hinder reliable operation. To address this challenge, this paper delves into a LiDAR-based approach for SLAM. LiDAR offers substantial advantages, including a marked reduction in susceptibility to variations in lighting conditions and the ability to operate at high frequencies. However, established 2D LiDAR-based mapping methods, such as GMapping and Hector SLAM, showcase success primarily in controlled indoor environments [9]. These methods struggle to adapt to environments characterized by sparsity, featuring long corridors, tunnels, or open spaces.
To address these limitations of single sensors, recent research has explored multi-sensor fusion techniques by leveraging the strengths of complementary sensors. LIO-SAM [10] integrates an IMU not only to predict robot motion prior to LiDAR scan matching, but also to initiate the LiDAR odometry optimization process. This combined approach refines IMU sensor error estimates within a factor graph. Wang et al. [11] proposed a multimodal system that integrates LiDAR, camera, IMU, encoder, and GPS to estimate the robust and accurate state of the robot. It can also eliminate dynamic targets and unstable features. Overall, tightly coupled multi-sensor fusion systems are becoming essential for ensuring the accuracy and flexibility required for autonomous robots operating in complex environments.
@@ 4 Conclusion
This paper presented an effective multisensor fusion approach for SLAM in textureless environments. Our method integrates a sensor selection and weighting strategy that leverages sensor characteristics to prioritize high-quality measurements. This combined approach significantly improves the accuracy of robot positioning and mapping. The factor graph framework enables flexible adaptation and integration of various sensor types within the system. Our experimental evaluation demonstrated the ability of our method compared to Karto SLAM, achieving substantial reductions in both positioning error (67.68%) and angular error (26.98%). Future work will focus on developing a mechanism for loop closure, which is crucial for global consistency in SLAM. Additionally, we plan to conduct real-world experiments to validate the effectiveness of our approach in practical scenarios.

### v2

@@ 1 Introduction
The field of Autonomous Mobile Robots (AMRs) is experiencing a surge of development across diverse industries such as healthcare, security monitoring [1][2], and construction [3]. To thrive in these dynamic environments and fulfill their designated tasks, AMRs necessitate robust capabilities in three critical areas: information gathering [4], environment perception [5], and precise Simultaneous Localization and Mapping (SLAM) [6][7][8]. Although various research efforts have explored effective mapping techniques that utilize individual sensors such as LiDAR [9] or cameras [10], cameras have significant limitations. Their sensitivity to light fluctuations can considerably hinder reliable operation. To address this challenge, this paper delves into a LiDAR-based approach for SLAM. LiDAR offers substantial advantages, including a marked reduction in susceptibility to variations in lighting conditions and the ability to operate at high frequencies. However, established 2D LiDAR-based mapping methods, such as GMapping and Hector SLAM, showcase success primarily in controlled indoor environments [11]. These methods struggle to adapt to environments characterized by sparsity, featuring long corridors, tunnels, or open spaces.
To address these limitations of single sensors, recent research has explored multi-sensor fusion techniques by leveraging the strengths of complementary sensors. LIO-SAM [12] integrates an IMU not only to predict robot motion prior to LiDAR scan matching, but also to initiate the LiDAR odometry optimization process. This combined approach refines IMU sensor error estimates within a factor graph. Several previous works [13][14] proposed a multimodal system that integrates LiDAR, camera, IMU, encoder, and GPS to estimate the robust and accurate state of the robot. It can also eliminate dynamic targets and unstable features. Overall, tightly coupled multi-sensor fusion systems are becoming essential for ensuring the accuracy and flexibility required for autonomous robots operating in complex environments.
@@ 4 Conclusion
This paper presented an effective multisensor fusion approach for SLAM in textureless environments. Our method integrates a sensor selection and weighting strategy that leverages sensor characteristics to prioritize high-quality measurements. This combined approach significantly improves the accuracy of robot positioning and mapping. The factor graph framework enables flexible adaptation and integration of various sensor types within the system. Our experimental evaluation demonstrated the ability of our method compared to Karto SLAM, achieving substantial reductions in both the positioning error (67. 68%) and the angular error (26. 98%). Future work will focus on developing a loop closure mechanism, which is crucial for global consistency in SLAM. In addition, we plan to conduct real-world experiments to validate the effectiveness of our approach in practical scenarios.